# WisPerMed at BioLaySumm: Adapting Autoregressive Large Language Models for Lay Summarization of Scientific Articles

### v1

@@ 2 Dataset
The dataset (Goldsack et al., 2022) of the Shared Task (Goldsack et al., 2024) contains two collections of scientific journal articles and the corresponding lay summaries, namely PLOS and eLife. Lay summaries of the PLOS dataset were written by the authors of the articles, while eLife lay summaries were written by expert editors in correspondence with the authors.
For the remainder of this paper, any reference to the validation set or test set will include eLife and PLOS unless otherwise specified.
@@ 3 Evaluation Metrics
The generated summaries were evaluated using ten metrics categorized under relevance, readability, and factuality. Relevance was measured with Recall-Oriented Understudy for Gisting Evaluation (Lin, 2004) (ROUGE-1, ROUGE-2, ROUGE-L) and BERTScore (Zhang et al., 2020). Readability was evaluated using the Flesch-Kincaid Grade Level (FKGL) (Kincaid, 1975), Dale-Chall Readability Score (DCRS) (Chall and Dale, 1995), Coleman-Liau Index (CLI) (Coleman and Liau, 1975), and Learnable Evaluation Metric for Text Simplification (LENS) (Maddela et al., 2023). Factuality was assessed with AlignScore (Zha et al., 2023) and Summary Consistency (SummaC) (Laban et al., 2022).
@@ 4.2 Prompt Variations
Prompts can guide the LLM’s content generation process without the need for fine-tuning. In the zero- and few-shot settings, different prompt variations and their effect on the evaluation metrics were examined. In the few-shot setting, examples from the training and validation set were included in the prompt when performing inference on the validation and test set, respectively. The format of the few-shot prompts is designed to emulate a preceding conversation with the model, with the examples serving as answers from the model.
To choose the best few-shot examples, all examples were ranked based on their average readability and factuality. The two and three highest-ranked examples were selected for the eLife and PLOS datasets, respectively.
@@
Three prompt variations were created which provide the model with different kinds of context information. It was decided that BioM would be utilized for all experiments involving these variations due to its superior performance on the validation set in the few-shot setting (see Table 3 in Appendix D). The first prompt variation includes a persona description of a science communicator (BioMpers). The model is then instructed to channel the expertise of the described persona to craft the lay summary based on the abstract. The second prompt variation is a modification of the initial prompt, but it includes the introduction as further context for background information (BioMintro). The third prompt variation contains the abstract and a guide on how to write a lay summary, accompanied by instructions concerning the content and style of the requested lay summary (BioMguide). The wording of all prompts can be found in Appendix A.
@@ 4.3 Dynamic Expert Selection (DES)
The success of a model depends on factors (Ling et al., 2024) such as the nature of the data, the complexity of the domain, and the design of the prompt. Consequently, a model may yield a more suitable lay summary when prompted in a different manner. In consideration of this assumption, a Dynamic Expert Selection (DES) was developed. It selects one text from a set of candidate texts based on metrics that do not necessitate the target lay summary as a reference. The mechanism uses the readability metrics FKGL, DCRS, and CLI, as well as the factuality metrics. These metrics are computed for each candidate text and a min-max normalization is applied to each score so that the values are between 0 and 1. Prior to the normalization the readability metrics were multiplied by -1 so that 1 is the best and 0 is the worst. After computing the mean of all readability and factuality scores, the overall score is computed. Since the target lay summaries in eLife have a higher readability than those in PLOS (Goldsack et al., 2022), the overall scores are computed with different weights for the two aspects. For eLife candidates, the weights are set to 0.675 and 0.325, whereas for PLOS candidates, the weights are set to 0.25 and 0.75 for readability and factuality, respectively.
@@ 5 Results
BioM exhibits the highest LENS score in the zero-shot setting. However, its performance in relevance and factuality is the lowest. Few-shot learning resulted in enhanced performance across all metrics except for LENS. The persona prompt (BioMpers) led to an improvement in relevance. Including the introduction in the prompt (BioMintro) resulted in a reduction in all aspects despite the fact that the model had access to more information from the article itself. In comparison, the prompt with the guide (BioMguide) exhibits minimal enhancements. The optimal few-shot learning for BioM occurred with the initial prompt, which achieved the highest readability and factuality in the few-shot setting, excluding the DES approach. However, OpenBio slightly underperformed with this prompt in the few-shot setting, except for the LENS score, where it performed best in this setting.
@@ 6 Conclusion
This paper presents the WisPerMed team’s approaches to automatic lay summarization within the biomedical domain, utilizing a combination of fine-tuning, prompt variations, and DES.
Among these approaches, fine-tuning emerged as an effective method, leading to the best performance across most metrics. This underscores the importance of task-specific training in optimizing model output for complex summarization tasks. Additionally, BioM demonstrated strong few-shot learning capabilities, illustrating its robustness and versatility in generating accurate and relevant summaries even without extensive training. As the model adjusts to the factuality and readability of given examples, providing better examples could lead to further enhancements in these aspects.
BioM demonstrated high factuality, even when provided solely with abstracts as input, suggesting that BioM leveraged domain-specific knowledge acquired during pre-training. This indicates that domain adaptation remains an important factor when using LLMs for lay summarization of scientific articles, as BioM outperformed the larger general model Llama3. A potential limitation of this study is the possibility that BioM may have been previously exposed to the gold standard.
The DES mechanism refined readability and factuality by retrospectively selecting the best text outputs based on evaluation metrics. This highlights the potential of metric-driven selection to improve the quality of lay summaries further.

### v2

@@ 2 Dataset
The dataset (Goldsack et al., 2022) of the Shared Task (Goldsack et al., 2024) contains two collections of scientific journal articles and the corresponding lay summaries, namely PLOS and eLife. PLOS and eLife also include the section headings and keywords of the article. The PLOS dataset has 24,773 examples in the training split and 1,376 examples in the validation split, whereas the eLife dataset is smaller with 4,346 examples in the training split and 241 examples in the validation split. The test split consists of 142 examples for both datasets. Lay summaries of the PLOS dataset were written by the authors of the articles and are approximately 150-200 words long, while eLife lay summaries were written by expert editors in correspondence with the authors and are about twice as long.
For the remainder of this paper, any reference to the validation or test set will include eLife and PLOS unless otherwise specified.
@@ 3 Evaluation Metrics
The generated summaries were evaluated across ten metrics that fall into the following categories: relevance, readability, and factuality. Relevance was assessed through Recall-Oriented Understudy for Gisting Evaluation (Lin, 2004) (ROUGE-1, ROUGE-2, ROUGE-L) and BERTScore (Zhang et al., 2020). ROUGE counts the overlapping n-grams in the generated texts and target lay summaries, whereas BERTScore uses contextual word embeddings to compare the semantic similarity of the two texts. Readability was evaluated using the Flesch-Kincaid Grade Level (FKGL) (Kincaid, 1975), Dale-Chall Readability Score (DCRS) (Chall and Dale, 1995), Coleman-Liau Index (CLI) (Coleman and Liau, 1975), and Learnable Evaluation Metric for Text Simplification (LENS) (Maddela et al., 2023). The FKGL uses sentence lengths and syllable count per word to estimate readability. The DCRS uses a word list to compute the occurrences of words unknown to most 4-th graders and the CLI estimates the grade level necessary to comprehend the text. The LENS metric is a learnable evaluation metric trained on datasets containing human ratings of simplifications. In this setting, LENS measures the simplification of the abstract by the generated text using the target lay summary as a reference. Factuality was assessed with AlignScore (Zha et al., 2023) and Summary Consistency (SummaC) (Laban et al., 2022). The AlignScore quantifies the degree of alignment between the facts in the summary and the scientific article, while SummaC also includes consistency.
@@ 4 Methods and Experiments
This section outlines the methodology employed in the experiments conducted on the specified dataset.
@@ 4.2 Prompt Variations
Prompts can guide the LLM’s content generation process without the need for fine-tuning. In the zero- and few-shot settings, different prompt variations and their effect on the evaluation metrics were examined. In the few-shot setting, example lay summaries from the training and validation set were included in the prompt when performing inference on the validation and test set, respectively. The format of these few-shot prompts is designed to emulate a preceding conversation with the model, with the included examples serving as the model’s previous responses.
To choose the best few-shot examples, all examples were ranked based on their average normalized readability and factuality. The two and three highest-ranked examples were selected for the eLife and PLOS datasets, respectively.
@@
Additionally, three prompt variations were created, which provide the model with different kinds of context information. It was decided that BioM would be utilized for all experiments involving these variations due to its superior performance on the validation set in the few-shot setting (see Table 3 in Appendix D). LLMs can assume different roles and adapt their vocabulary accordingly (Salewski et al., 2023), resulting in enhanced performance in tasks related to the specified role. Accordingly, the first prompt variation comprises a persona description of a science communicator (BioMpers), instructing the model to utilize the expertise of this persona to create the lay summary based on the abstract. The model is then instructed to channel the expertise of the described persona to craft the lay summary based on the abstract. The second prompt variation is a modification of the initial prompt, incorporating the introduction to provide additional background information because associated context can improve LLM performance (Karmaker and Feng, 2023). The second prompt variation is a modification of the initial prompt, but it includes the introduction as further context for background information (BioMintro). The third prompt variation includes the abstract and a guide on how to write a lay summary (BioMguide), accompanied by instructions concerning the content and style of the requested summary. This method leverages the importance of clear and detailed task directives. The selection of these prompts was based on a few preliminary experiments with the model and an initial assessment of the responses. However, no comprehensive optimization was performed. The wording of all prompts can be found in Appendix A.
@@ 4.3 Dynamic Expert Selection (DES)
The success of an LLM depends on factors such as the properties of the dataset, the complexity of the domain, and the design of the prompt (Ling et al., 2024). Consequently, a model may yield a more suitable lay summary when prompted in a different manner. In addition, the output quality depends upon the selection of the inference parameters (Minaee et al., 2024). In consideration of this assumption, a Dynamic Expert Selection (DES) was developed. It selects the most appropriate text from a set of candidate texts based on metrics that do not require a reference lay summary.
The mechanism uses the readability metrics FKGL, DCRS, and CLI, as well as the factuality metrics AlignScore and SummaC. These metrics are computed for each candidate text. Readability scores are multiplied by -1 so that higher scores indicate better readability. All scores are normalized using min-max normalization to range between 0 and 1, where 1 is the best and 0 is the worst. For each candidate text, an overall score is calculated by multiplying the means with different weights. Given that the target lay summaries in eLife have a higher readability than those in PLOS (Goldsack et al., 2022), the overall scores are computed with different weights for the two aspects. For eLife summaries: Readability is weighted at 0.675 and factuality at 0.325. For PLOS summaries: Readability is weighted at 0.25 and factuality at 0.75. The candidate text with the highest overall score is selected as the most suitable lay summary. The selection of the weights is based on the assumptions about the target texts and comparisons of the overall scores on the validation dataset.
@@ 5 Results
BioM exhibits the highest LENS score in the zero-shot setting. However, its relevance and factuality performance are the lowest. Few-shot learning resulted in enhanced performance across all metrics except for LENS. The persona prompt (BioMpers) led to an improvement in relevance. Including the introduction in the prompt (BioMintro) resulted in a reduction in all aspects despite the fact that the model had access to more information from the article itself. In comparison, the prompt with the guide (BioMguide) exhibits minimal enhancements. The optimal few-shot learning for BioM occurred with the initial prompt, which achieved the highest readability and factuality in the few-shot setting, excluding the DES approach. However, OpenBio slightly underperformed with this prompt in the few-shot setting, except for the LENS score, where it performed best in this setting.
@@ 6 Conclusion
This paper presents the WisPerMed team’s approaches to automatic lay summarization within the biomedical domain, utilizing a combination of fine-tuning, prompt variations, and Dynamic Expert Selection.
Among these approaches, fine-tuning emerged as an effective method, leading to the best performance across most metrics. This underscores the importance of task-specific training in optimizing model output for complex summarization tasks. Additionally, BioM showed strong few-shot learning capabilities, illustrating its robustness and versatility in generating accurate and relevant summaries even without extensive training. As the model adjusts to the factuality and readability of given examples, providing better examples could lead to further enhancements in these aspects.
BioM reached high factuality, even when provided solely with abstracts as input, suggesting that BioM leveraged domain-specific knowledge acquired during pre-training. This indicates that domain adaptation remains an important factor when using LLMs for lay summarization of scientific articles, as BioM outperformed the larger general model Llama3.
The four prompt variations exhibited differing effects on the evaluation metrics. BioM is adept in fulfilling the role of a science communicator (BioMpers), as evidenced by the enhanced relevance. BioMintro and BioMguide did not significantly enhance the metrics, indicating that the increase in context was not beneficial for all texts. Without DES, a shorter prompt (BioMinitial) yielded the optimal results, suggesting that the model effectively comprehends the task from the provided examples. The DES mechanism further refined readability and, in particular, factuality by retrospectively selecting the best text outputs based on evaluation metrics. This highlights the potential of metric-driven selection to improve the quality of lay summaries further.
@@
Limitations
Only four discrete prompts in combination were tested with DES, and only two sets of inference parameters were explored. This limited scope means that the findings may not fully capture the potential variability and performance of the the various models under different conditions. The weights for the Dynamic Expert Selection method were chosen based on heuristics without any formal optimization, which could impact the robustness and generalizability of the results. Another limitation is the possibility that BioM may have been previously exposed to the gold standard summaries. If this is the case, it could skew the results by artificially inflating the model’s performance. These limitations indicate potential avenues for future research, including the necessity for more comprehensive prompt engineering, optimization of DES weights, and a wider range of tasks to ensure the robustness of the approach. Another potential future direction is adapting these methods for other complex domains or languages and exploring additional metrics.
Acknowledgement